Compound F has a molecular formula of C9H8O. What is the unsaturation number; Functional groups that are ppossibly present in compound F?

I keep getting this question wrong, can anyone help me?

7 of 29 Incorrect Question 27 of Map Compound F has a molecular formula of CgHgO. What is the Unsaturation Number (UN) or Double-Bond Equivalent (DBE)? Number Examine the IR spectrum provided for Compound F Based on the molecular formula, the UN/DBE you calculated, and the IR spectrum, check all the functional groups that are possibly present in compound F? carboxylic acid nitrile ketone alcohol □ alkyne benzene ring ring alkene amine Incorrect. aldehyde ester
Answer #1

Sol:-

Double bond equivalents (DBE) = C - X/2 - H/2 + N/2 + 1

Double bond equivalents (DBE) = 9 - 8 / 2 + 1

Double bond equivalents (DBE) = 9- 4 +1

Double bond equivalents (DBE) = 6

The functional groups are :-

Ring, benzene ring, aldehyde, ketone , alkene, alkyne